OpenStudy (anonymous):

show that, f(x) = x^3 is decreasing on [-1,1]

OpenStudy (anonymous):

Sorry, x^3 is an increasing function for all x! Reason is because the derivative of x^3 is 3x^2 and 3x^2 is positive when x> 0 and when x <0.

OpenStudy (anonymous):

Where did you get that question from?

OpenStudy (anonymous):

if it is x^3+(1/x^3) then

OpenStudy (anonymous):

Now, you are asking a new question. Use what you have learned to determine the answer to your next question.

OpenStudy (anonymous):

i first calcualte it s derivative vat after that

OpenStudy (anonymous):

set derivative equal to zero or where it is undefined . That will give you any critical points. Set up your chart..and where f'(x) > 0, the graph is increasing, and where f'(x) < 0, the graph is decreasing.

OpenStudy (anonymous):

That's the First Derivative Test.

OpenStudy (anonymous):

just give me help how to start it with

OpenStudy (anonymous):

OK. Watch.

OpenStudy (anonymous):

f(x) = x^3 + 1/x^3 First thing I would do is to rewrite the function as F9x) = x^3 + x^(-3) becuase it is easier to take the derivative. Note that x^-3 = 1/x^3 So the derivative of x^3 + x^(-3), f ' (x) = 3x^2 -3x^(-4)

OpenStudy (anonymous):

So far, so good?

OpenStudy (anonymous):

till here,i have already done

OpenStudy (anonymous):

Now, we want to take the derivative and find out where does f ' (x) = 0 and where is f ' (x) undefined.

OpenStudy (anonymous):

The best way to solve, may be to rewrite the derivative as 3x^2 - 3/x^4 And that is (3x^6 - 3)/x^4 Now, the fraction is zero when the numerator is zero: 3(x^6 - 1) = 0, so x = 1 and x = -1 is the solution. The derivative is undefined when x = 0 (becuase the denominator will be zero). So there are 3 critical points: x = -1, 0, 1.

OpenStudy (anonymous):

Now we look at four intervals: where x < -1 where -1 < x <0 where 0 < x < 1 where x > 1 Take a number in each interval and detrmine if f ' (x) is positive or negative. Example, in the first interval, where x < -1, take x = -2..if you plug in x = -2 into f ' (x) you get a positive number. Do the same for all 4 intervals. Where the derivative is positive, the function is increasing. So we know that when x < -1, the function is increasing. So work out the rest of the problem for each of the three remaining intervals.

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

fine.

OpenStudy (anonymous):

Want to know another way to do it?

OpenStudy (anonymous):

@msingh Do you?

OpenStudy (anonymous):

yes, if there is another way i will also go for it.

OpenStudy (anonymous):

By definition. \(f(x)\) is increasing on an interval \([a,b]\) if \(\forall x_1,x_2\in [a,b]\):\[ x_1<x_2 \implies f(x_1)<f(x_2) \] Start with: \[ x_1<x_2 \]Now consider how \((x_1)^2\) is positive as well as \((x_2)^2\), since they are squared. There are two cases. Case 1: suppose \((x_1)^2<(x_2)^2\): \[ (x_1)(x_1)^2<(x_2)(x_1)^2<(x_2)(x_2)^2\implies (x_1)^3<(x_2)^3\implies f(x_1)<f(x_2) \] Case 2: suppose \((x_2)^2<(x_1)^2\): This is only possible if \(x_1\) is negative, and will flip the inequality. Multiply by \(x_1\) then gives: \[ x_2(x_2)^2>x_1(x_2)^2>x_1(x_1)^2\implies f(x_2)>f(x_1)\implies f(x_1)<f(x_2) \] So for all casese:\[ x_1<x_2 \implies f(x_1)<f(x_2) \]

OpenStudy (anonymous):

I think the question should be changed to: show that, f(x) = x^3 is increasing on [-1,1]

OpenStudy (anonymous):

It's actually increasing everywhere.

OpenStudy (anonymous):

sry bro, qquestion was pt, f(x) = x^3 +( 1/x^3 ) is decreasing on [ -1, 1 ]

OpenStudy (anonymous):

I showed you this an hour ago.